OpenStudy (anonymous):

I really need help... I need to find a equation for this word sentence... You have been asked to organized the Homecoming Dance. you rented "The Gym" for $5000. "The Gym" has a capacity of 700 people. You decide to charge EACH person $25.00 admission. How many people will need to sign up and play if you want to cover your costs? I need to put this into a table and then into a graph I know how to do that though I just don't know how to put it into an equation... hELP please!!!!

OpenStudy (anonymous):

Please help...

OpenStudy (anonymous):

Walk me step by step my teacher didn't really teach me

OpenStudy (anonymous):

Just use a linear function.

OpenStudy (anonymous):

What's that?

OpenStudy (anonymous):

A function which describes a line. It looks like this: \[f(x) = ax + b\]

OpenStudy (anonymous):

Okay so how do I plug in the numbers into that formula??

OpenStudy (anonymous):

:( I just guess or something... Goodnight I seriously hate this website like seriously wow...this is worse then my ex bf...

jimthompson5910 (jim_thompson5910):

Let x = number of people who attend the dance

jimthompson5910 (jim_thompson5910):

if it costs $25 a person, and x people pay this price, then this will bring in 25x dollars

jimthompson5910 (jim_thompson5910):

it costs $5000 to rent out the gym

jimthompson5910 (jim_thompson5910):

So the profit is given by Profit = Revenue - Cost P = R - C P = (25x) - (5000) P = 25x - 5000 To figure out how many students must attend to break even, you plug in P = 0 and solve for x P = 25x - 5000 0 = 25x - 5000 5000 = 25x 25x = 5000 x = 5000/25 x = 200 So 200 people need to show up and pay $25 each to break even

jimthompson5910 (jim_thompson5910):

To graph P = 25x - 5000, which is really the same as y = 25x - 5000, you just need two points If x = 0, then y = 25x - 5000 y = 25(0) - 5000 y = 0 - 5000 y = -5000 Which means that the point (0,-5000) is on the line. You can add this to the table as well. If x = 50, then y = 25x - 5000 y = 25(50) - 5000 y = 1250 - 5000 y = -3750 Which means that the point (50,-3750) is on the line. You can add this to the table as well. You have the two points (0,-5000) and (50,-3750). Plot them on an xy axis and draw a straight line through them to graph the line